34) If three 2.4-uF capacitors are connected in series, what is the combined capacitance? A) 8.0E-7 F B) 1.4E-5 F C)7.2E-8 F D) 2.4E-6F E) 4.6E-6 F
Answer #1

In Series combination. 1/Ceq = 1/C1 +1/C2+1/C3

Here, Three capacitors of equal magnitude is connected in series.

So, Ceq = C/3 = (2.4 x 10-6)/3 = 8.0 x 10-7 F.

So, Ans will be (A) 8.0E-7 F